Alex spent 3/7 of his money. He gave 1/4 of the remainder to his sister. He had $120 left. How much money did he have in the beginning?
Answer by rfer(16322)

120/(3/4)=
120*4/3=160
----------
160/(4/7)=
160*7/4=$280